### v1

# Augmenting NER Datasets with LLMs: Towards Automated and Refined Annotation

Abstract
In the field of Natural Language Processing (NLP), Named Entity Recognition (NER) is recognized as a critical technology, employed across a wide array of applications. Traditional methodologies for annotating datasets for NER models are challenged by high costs and variations in dataset quality. This research introduces a novel hybrid annotation approach that synergizes human effort with the capabilities of Large Language Models (LLMs). This approach not only aims to ameliorate the noise inherent in manual annotations, such as omissions, thereby enhancing the performance of NER models, but also achieves this in a cost-effective manner. Additionally, by employing a label mixing strategy, it addresses the issue of class imbalance encountered in LLM-based annotations. Through an analysis across multiple datasets, this method has been consistently shown to provide superior performance compared to traditional annotation methods, even under constrained budget conditions. This study illuminates the potential of leveraging LLMs to improve dataset quality, introduces a novel technique to mitigate class imbalances, and demonstrates the feasibility of achieving high-performance NER in a cost-effective way.
1 Introduction
In the field of Natural Language Processing (NLP), Named Entity Recognition (NER) is a critical task that involves identifying and classifying named entities in a given text into predefined categories, including persons, organizations, locations, temporal expressions, quantities, monetary values, and percentages (Mohit, 2014). NER is integral to a wide range of applications, from basic information retrieval and content classification to more complex tasks like question answering and machine translation (Guo et al., 2009; Aliod et al., 2006; Babych & Hartley, 2003). The enhancement of NER model performance not only directly contributes to the improvement of broad NLP applications but also plays a crucial role in enhancing our understanding and engagement with text-based data (Cheng & Erk, 2019; Etzioni et al., 2005).
In recent years, NER methods leveraging the in-context learning capabilities of Large Language Models (LLMs) have been proposed, demonstrating high performance close to that of models specialized in NER by carefully selecting the examples provided to the LLMs (Wang et al., 2023). This indicates the effectiveness of utilizing LLMs’ advanced text comprehension abilities for NER tasks. On the other hand, inference with LLMs requires significant computational resources and involves lengthy latency before obtaining inference results, making them unsuitable for applications requiring the rapid processing of huge amounts of text data. Therefore, we believe it is crucial to develop approaches that leverage the language understanding capabilities of LLMs while requiring fewer computational resources and achieving faster NER task processing. Consequently, we focus on improving the quality of datasets used for training NER models with the aid of LLMs, aiming to enhance the performance of smaller, faster, dedicated NER models.
The performance and effectiveness of NER models rely heavily on the quality of annotations in the datasets used during their training (Mitkov et al., 2000). Currently, the majority of datasets employed for training NER models are annotated manually by humans. This manual annotation process is fraught with several challenges, such as inconsistent quality of annotations, the presence of missed and incorrect annotations, significant time and financial costs, and the need for ethical considerations when dealing with text data that may contain offensive content, among others (Katiyar & Cardie, 2018; Fort et al., 2011).
In this study, we propose a novel approach to address these issues, particularly the problems of missed annotations and the time and financial costs associated with manual annotation, intending to enhance the quality of datasets and thereby strengthen the performance and effectiveness of NER models. Our proposed method leverages LLMs to supplement the missed annotations in datasets that have been manually annotated (where the presence of missed annotations is assumed) as shown in Figure 1. By utilizing a hybrid approach that combines manual annotations with annotations generated by large LLMs, we aim to recover from the noise introduced by missed annotations in the dataset, thereby enhancing the quality of the datasets used for training NER models at a low cost and in an automated manner.
Figure 1: LLM-based annotation and label mixing.
In our proposed method, a detailed analysis of the annotations by LLMs revealed that a single expression could be assigned multiple named entity labels. Also, an imbalance in the data volume per label exists, which in turn causes performance disparities among labels in the NER model. This is particularly critical in the “Miscellaneous” category, where such inequalities are most pronounced. To address these issues, we implemented a label mixing technique for the multiple assigned named entity labels, thereby ensuring the robustness of the NER model by mixing these labels.
In the experiments, the CoNLL03 (Tjong Kim Sang & De Meulder, 2003) and WikiGold (Balasuriya et al., 2009) datasets were enhanced using our proposed hybrid annotation method combining manual human efforts and LLMs, and the improved datasets were then used to train a BERT-based (Devlin et al., 2019) NER specialized model. The experimental results demonstrated that our proposed approach could achieve high performance in training NER models with datasets containing noise such as missed annotations, despite requiring minimal financial cost.
The contributions of this study can be summarized in three main points:
  • We propose a hybrid annotation method combining manual human efforts and LLMs for NER datasets, automatically recovering from noise such as missed annotations contained within the datasets (Figure 2).
  • By incurring only a minimal financial cost, we enhance the performance of NER-specialized models (Table 1).
  • In response to the issue of imbalance among named entity labels arising from LLM-based annotations, we propose a label mixing technique, ensuring the robustness of the NER model (Figure 4).
2 Related Works
2.1 Reliability of NER datasets
The reliability of datasets used in NER tasks has been scrutinized. Studies, such as the one presented in Han et al. (2023a), highlight the inherent errors in these datasets. These inaccuracies range from mislabeled entities to inconsistent categorization, impacting the overall performance and reliability of NER systems (Mishra et al., 2020; Mitkov et al., 2000).
2.2 LLMs for NER tasks
Wang et al. (2023) attempted to utilize GPT, a large-scale language model (LLM) (Brown et al., 2020; Ouyang et al., 2022), for annotation tasks and validated its effectiveness. Additionally, they tested the output format and improved annotation accuracy using few-shot learning. LLMs’ advantages for NER tasks are high consistency and efficiency. Conventional methods trained with manual annotation tend to vary due to annotator subjectivity and fatigue. In addition, the annotation cost for LLMs is much cheaper than that of human annotators. The disadvantage of their work is that the computational cost of inference is very high each time. While their approach directly solves the NER task, which differs, our approach includes augmenting and creating the NER dataset from scratch.
2.3 Challenges and Remedies in Automatic Annotation
There are several challenges in automatic annotation. First, Ghosh et al. (2023) proposes ACLM (Attention-map aware keyword selection for Conditional Language Model fine-tuning), an attention map-based data expansion method to address the problem of data imbalance. It should be noted, however, that this method does not take advantage of LLMs. Next, Shen et al. (2023b) proposes using multiple prompt templates to address the problem of multiple label assignments to the same entity. Furthermore, DiffusionNER in Shen et al. (2023a) and Zhang et al. (2022) address the issues of noise and generative bias by utilizing causality-based methods. To address these issues, we propose a hybrid approach that combines manual annotation with automatic LLMs annotation.
Other related studies in NER that must be mentioned are Li et al. (2020) for a comprehensive review of NER deep learning methods, Liu et al. (2021) addresses noise from external knowledge bases, and Wang et al. (2022) propose MINER (Mutual Information based Named Entity Recognition) for out-of-vocabulary support.
While the above diverse approaches address the NER challenge, this study proposes a unique methodology that focuses on improving the reliability of the dataset itself and leads to better accuracy for NER tasks.
3 Methods
LLM-based Annotation
In linguistic data automation using LLM, the GPT-NER methodology, as delineated in Wang et al. (2023), emerges as a paradigmatic example. This method profoundly influences our study, harnessing LLMs to facilitate automated annotation. Central to the GPT-NER strategy is the LLM’s capability to discern and earmark specific entities within the text, encapsulating these segments with distinct character strings, specifically “@@” and “##”. For illustrative purposes, given an input such as “I live in Tokyo”, wherein the objective is to pinpoint LOC (location) entities, the LLM’s output would be “I live in @@Tokyo##”. This delineation effectively signifies the accurate detection of “Tokyo” as a location entity, with the “@@” and “##” serving as the entity’s boundaries.
Additionally, this approach empowers the LLM to apprehend both the intrinsic properties of entities and their interplay with the input-output schema, achieved through the integration of few-shot examples. In alignment with this methodology, we have formulated the ensuing prompt structure. {itembox}[l]Instructions for the Large Language Model You are an excellent linguist.
Identify [target entity type] entities in given text.
If the text contains no [target entity type] entities, replicate the input text without any changes.
Strictly adhere to the definition provided below.
Definition of [target entity type]:
[definition target entity type]
[Example $1$ Input]
[Example $1$ Output]
[Example $2$ Input]
[Example $2$ Output]
$\cdots$
[Example $K$ Input]
[Example $K$ Output]
Within this structured prompt design, the [target entity type] slot accommodates various entity types, including but not limited to PER, ORG, LOC, and MISC. Concurrently, the [definition target entity type] slot is allocated for the explicit definition of these enumerated entity types. Moreover, the [Example $K$ Input] and [Example $K$ Output] slots are respectively reserved for the input text and its corresponding output text, collectively functioning as few-shot examples.
As this prompt dictates, the entity recognition process is methodically applied to each target entity type within the text targeted for annotation. In scenarios where this strategy involves multiple iterations of annotation for each type of entity, a single word may be assigned multiple different entity types. To resolve instances of such overlapping entities, the simplest method we adopt is to select one entity from those multiple assignments randomly.
The criterion for selecting few-shot examples is inspired by the GPT-NER approach, entailing the selection of examples from validation data. This selection is predicated on the cosine similarity between token embeddings present in the input text.
Hybrid: LLM-Based and Manual Annotation
To enhance the quality of the NER dataset, we created Hybrid annotations by merging annotations from LLMs with those generated manually. This merging process involved carefully analyzing the annotations from both sources, ensuring consistency and accuracy. The merged data combines all entity labels from the manual data and those from LLM annotation that do not overlap the same position as those from the manual data. Since the entity labels of the manual data are expected to be of higher quality than the ones from LLM annotation, they are merged in preference to the ones from LLM annotation. The merged dataset offers a more robust and diverse set of annotations, mitigating biases and errors that might exist in either of the individual sources.
Figure 2: Comparison of model performance when using noised and datasets annotated by Hybrid approach.
Label Mixing in LLM-Based Annotation
Addressing the issue, as mentioned earlier, of multiple labels being assigned to a single word, we extended the Mixup (Zhang et al., 2018) technique, traditionally used in image processing, to NER tasks. Mixup is a data augmentation method that improves model performance and robustness by mixing two inputs and outputs. Our approach involves blending two different entity labels for a single token to create new annotations. This method is formally represented as follows:
$y_{\text{mix}}=\lambda\times y_{i}+(1-\lambda)\times y_{j}$
Where $y_{i}$ and $y_{j}$ are two labels annotated NER tags, and $\lambda\in[0,1]$ is a parameter determining the mixing ratio. Following the mixup technique (Zhang et al., 2018), $\lambda$ is drawn from the beta distribution ($\alpha=\beta=0.2$). Compared to randomly selecting one label from multiple labels assigned to a single word, this technique contributes to a more nuanced understanding of tokens that could belong to multiple entity types, enhancing the dataset’s richness and classification flexibility.
4 Experiments
Budget #Manual #LLM F1 Score
\$38 87 0 0.02
70 702 0.81
35 2106 0.84
0 3510 0.85
\$152 351 0 0.63
280 2808 0.85
140 8424 0.87
0 14041 0.86
\$608 1404 0 0.86
1263 5616 0.87
1123 11232 0.87
Table 1: Comparison between different data count balances from manual and LLM-based data under the fixed budgets.
4.1 Setup
In our research, we conducted experiments using two distinct datasets. Firstly, we employed the CoNLL03 dataset (Tjong Kim Sang & De Meulder, 2003), recognized as a standard benchmark in the field of NER. Additionally, acknowledging recent advancements in this field models that have led to a saturation in performance improvements on the CoNLL03 dataset, we also included the WikiGold dataset (Balasuriya et al., 2009). The WikiGold dataset was specifically chosen due to its higher level of challenge and the observation that performance levels achieved on it have not yet paralleled those on the CoNLL03 dataset. This was split into training, validation, and test sets in a 7:1:2 ratio for the experiments.
For the automated annotation process, we leveraged ChatGPT (gpt-3.5-turbo-0613) API [^1] https://chat.openai.com as LLM, and we set the number of few-shot examples at 32.
Finally, to validate the effectiveness of our annotation approach, we fine-tuned a sequence classification model based on a pre-trained BERT (Devlin et al., 2019), utilizing the datasets annotated through this methodology. The evaluation of the NER task is in line with Tjong Kim Sang & De Meulder (2003), and we present weighted average F1 scores in this paper.
A detailed description of the hardware and software configurations is also provided in Appendix appendix A to ensure reproducibility.
4.2 Results
Noise Recovery with LLM-based Annotations:
We assessed the capability of LLM-based annotations to recover from noisy data. Han et al. (2023b) pointed out some annotation errors, such as missing annotations or label switching, in the CoNLL03 dataset (Tjong Kim Sang & De Meulder, 2003). We prepared noised data that reproduces this problem by removing annotations for some entity annotations. We show that it can be improved by combining LLM-based annotation in the previously described manner. We experimented with 20%, 40%, 60%, 80% entity annotations removed. By comparing the noised datasets to those enhanced through our proposed combining methodology, we established a baseline for improvement. Figure 2 presents a comparative analysis of these datasets, highlighting the efficiency of our approach to noise reduction. It was shown that even when noise is severe, it can be significantly improved by combining LLM-based annotations.
Figure 3: Number of annotations for each entity type in the manual, noised, and LLM-based datasets.
Comparison Under the Same Budget:
We evaluated our approach under identical budget constraints by drawing parallels with the study of  Ding et al. (2023). Our experiments were conducted on three different budgets. We created composite datasets including different numbers of samples from manual and LLM-based annotations. The middle budget (\$152) is the cost of LLM-based annotation of all training data of the CoNLL03 dataset, and the costs of a quarter and a quadruple of it are assumed for small (\$38) and large (\$608) budgets, respectively. The relationship between the balance of the number of data in each budget and performance is shown in Table 1.
When the budget is small (\$38), training is not possible with only manual data due to the small amount of data. Table 1 shows that it is thought that performance can be improved by using LLM-based annotations to increase the training data. For the middle budget (\$152), the model’s performance was reduced using only the manual data; using just 2,808 LLM-based annotations showed a significant improvement. In addition, the performance of the Hybrid annotations was higher than that of LLM-based alone, suggesting that the use of both annotations contributes to the performance. For the large budget (\$608), the performance is sufficient with 1,404 manual data, but a slight performance improvement was confirmed by reducing the manual data and using LLM-based as well. The results show that the composite datasets can increase the number of data at a low cost and improve the performance of downstream tasks under all budgets.
Addressing the Imbalanced Class Issue:
Figure 3 shows that there is a difference in the number of annotations for each entity type. This imbalance issue is even greater with annotation by LLM. Also, the performance difference of LLM-based annotation per entity type is large compared to manual annotation. To tackle the class imbalance issue, we motivated the use of our proposal, label mixing. We present the comparison between LLM-based annotation and label mixing in Figure 4. Label mixing reduced the performance difference between entity types. Therefore, we believe that soft labeling of LLM-based annotations according to possible entities rather than one-hot hard labels will prevent model training in a biased manner.
Figure 4: Improvement of each entity type and micro average with label mixing.
CoNLL03 WikiGold
ORG LOC MISC ORG LOC MISC
PER 72 37 15 13 5 1
ORG - 180 59 - 44 41
LOC - - 158 - - 5
Table 2: The number of entities overlapping among classes in the CoNLL03 and WikiGold dataset.
In addition to the analysis above, we further explore the impact of LLM-based annotations and label mixing on the CoNLL03 and WikiGold datasets, particularly focusing on the number of entities that overlap between two classes. Table 2 presents a detailed quantification of these overlapping entities between different entity types when annotated by LLM. The fact that the entities overlapping between classes constitute only 521 cases (2.65%) of the CoNLL03 dataset and 109 cases (6.80%) of the WikiGold dataset, and yet the application of label mixing to these specific data points results in a performance improvement across the dataset, is a suggestive observation. This observation implies that the strategic application of label mixing to even a small subset of data can improve the performance balance between entity types. Moreover, the fact that the number of entities overlapping between classes such as LOC and MISC with ORG is comparatively higher than other combinations, while there is less overlap between PER and MISC, indicates a potential challenge in the discriminatory capability of LLM-based annotations for the former combinations. These observations support the effectiveness of label mixing as a technique to buffer these shortcomings, suggesting that it serves as a valuable strategy for improving the robustness and performance of NER models.
5 Discussion and Conclusion
Our study extends the use of LLMs beyond annotation tasks. We identified and addressed new NER annotation challenges, proposing novel solutions. Our integration of LLM-based annotations with manual annotation, especially in datasets with noise or limited samples, significantly enhances downstream task performance. This method leverages LLMs’ evolving strengths (Zhao et al., 2023; Min et al., 2023) and introduces label mixing to address biases and imbalances from LLM-based annotations. Our work illuminates LLMs’ role in enhancing NER, underscoring the need for ongoing adaptation and improvement in annotation methodologies.
6 Ethical Considerations
In our research, we focus not on generating text, but on the automation of annotations. This approach places concerns about creating potentially harmful sentences outside our scope of study. However, while our primary objective is to annotate entities with labels, it is essential to consider the implications of certain annotations, such as the toxicity of entities, especially when these annotations pertain to specific demographic groups.
This risk stems from biases inherent in the LLMs used for annotation, necessitating continuous monitoring.
7 Limitations
Finally, we would like to mention the limitations of our work. As the first limitation, we have only validated our proposed methods with relatively small datasets, CoNLL03 and Wikigold, thus requiring validation with larger NER datasets, such as OntoNotes 5.0 (Weischedel et al., 2013).
Another limitation is that when using the ChatGPT (gpt-35-turbo-0613) API, some regions and organizations are not generated due to their alignment. It is conceivable that this kind of alignment-based distribution bias could affect NER task performance.
The final limitation involves the LLM-based annotation method for NER tasks, which requires computational costs proportional to the number of entity types. Thus, algorithms independent of entity type count are needed.
Appendix
Appendix A Details of Experiment Setup
Hardware Configurations:
In our experiments, we used virtual machines on Google Cloud. The detailed machine specifications are as follows:
  • Machine type: n1 standard 4
  • Memory: 15 GB
  • virtual CPU: Intel Broadwell x 4
  • GPU: NVIDIA T4 x 1
BERT Configurations:
The configurations of BERT used for the performance evaluation are as follows:
  • Pre-trained model: bert-model-unbased  [^2] https://huggingface.co/bert-base-uncased
  • Learning rate: $051e-$ 05
  • Upper limit for norm of gradients: $10$
  • Batch size: $64$ for training, $32$ for inference
  • Max length of token: $128$
Each training was terminated when the F1 score on the validation set remained unchanged for three consecutive epochs.
In NER, labels must be assigned at the word level, but BERT’s tokenizer often splits a single word into multiple tokens. Therefore, in this study, when a word is split into multiple tokens, when a word is split into multiple tokens, the model adopts the label for the first token as the label for the entire word.
Appendix B Datasets
CoNLL03 (Tjong Kim Sang & De Meulder, 2003):
The CoNLL-2003 dataset is designed to recognize four types of named entities, including person names, place names, organization names, and other miscellaneous entities. The data file consists of four columns: word, part-of-speech tag, syntactic chunk tag, and named entity tag, with each item separated by a space. The named entity and syntactic chunk tags use the I-TYPE format  [^3] This tag means ”Inside” and is assigned to words that are within an entity of a specific type (TYPE). That is, this tag is used when a word is part of a specific entity and indicates the type of that entity. For example, ”I-PER” indicates that the word is part of a person’s name (Person)., and the B-TYPE  [^4] The B in B-TYPE means ”Beginning” and is the tag given to the first word of the second entity when directly followed by an entity of the same type. This tag is used to indicate that a new entity is beginning and to distinguish it from the previous entity. For example, ”B-PER” means that the new person’s name begins with this word. tag is introduced to indicate the start of a new phrase.
WikiGold (Balasuriya et al., 2009):
The Wikigold dataset is a manually annotated corpus for named entity recognition consisting of a small sample of Wikipedia articles. Words are labeled for each of the four named entity classes (LOC, MISC, ORG, and PER) of CONLL-03 described above.